1 Review I Constants Public Radio station KXPR-FM in Sacramento broadcasts at 88.9 MHz. The radio waves pass between two tallThe answer isn't 2.08 or 2.58
Answer #1

a) Wavelength of the wave is, 3 x108 m/s 2= 88.9x10 Hz = 3.375 m Horizontal angle is, For m=1 m2 O= sin (1)(3.375 = sin 15 +1b) Intensity is, sin B 2 I=1. B 2 sin = 1, = πα sin θ 2 na sin e 2 sin T (15 m)sin (6.0) 3.375 (3.80 W/m A (15 m) sin(6.0) 3.